The density of an unknown element in the gaseous state is 1.60 g at 300 K and 1 atm. Which of the following could be the element?

a. He
b. Ne
c. Ar
d.O2
e. Cl2
Answer #1

Ideal gas equation PV = nRT

P = 1 atm

T = 300 K

Concentration = n/V = P/RT

= 1/(0.08206 x 300) = 0.04062 mol/L

Since density = 1.60 g/L

Mass of 0.04062 mol of element = 1.60 g

Molar mass = mass/moles

= 1.60/0.04062 = 39.4 g/mol

Thus the unknown element is c. Ar